OpenStudy (anonymous):

Solve on the interval [0,2π): A. 7π/6, 11π/6 B. π/6, 5π/6 C. π/3, 5π/3 D. 2π/3, 4π/3

OpenStudy (anonymous):

Kassia what should i solve

OpenStudy (anonymous):

could be sin x = -1/2 in which case A!! medal plz!!

OpenStudy (anonymous):

sorry, 4secx+6=-2

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

secx = -2.... cosx = -1/2 x = pi/3

OpenStudy (anonymous):

x = 2pi/3

OpenStudy (anonymous):

D will be good

OpenStudy (anonymous):

4/cosx = -8 cos x = -1/2 x = 2pi/3 , 4pi/3 D